The first two are answered correctly but the third is wrong Where am I going wrong?
Suppose that A and B are two events for which P(A) = 0.22, P(B) = 0.67, and P(B given A) = 0.51. Find each of the following:
A. P(A and B) =
.1122
B. P(A or B) =
.7778
C. P(A given B) =
.2 Incorrect...

Answer by Theo(13342)

p(a) = .22
p(b) = .67
p(b|a) = .51

p(b|a) = p(ab) / p(b)

you calculated for p(ab) and got .1122 which was right.

p(a|b) = p(ab) / p(b) = .1122 / .67 = .1674626866 which can be rounded to as many decimal digits as required.

round to 4 = .1675
round to 3 = .167
round to 2 = .17
etc.

try that.
see how it goes.